A school principal claims that grade 11 students have a mean grade of 86 with a standard deviation of 4 . Suppose that the distribution is approximately normal . What is the probability that a randomly selected grade will be greater than 82 but less than 90



1
Expert's answer
2021-05-21T07:31:31-0400

Let "X=" grade: "X\sim N(\mu,\sigma^2)"

Given "\mu=86, \sigma=4"

"P(82<X<90)=P(X<90)-P(X\leq82)""=P(Z<\dfrac{90-86}{4})-P(Z\leq\dfrac{82-86}{4})""=P(Z<1)-P(Z\leq-1)""\approx0.84134-0.15866\approx0.6827"